Start learning →

A and B are mutually exclusive, P(A)=0.3 and P(B)=0.45. Find P(A or B).

A0.75
B0.135
C0.15
D0.85
Answer & Solution
Correct answer: A. 0.75
For mutually exclusive events probabilities add: 0.3 + 0.45 = 0.75.
Practise this question with Quill — SSC CGL →

Keep practising Probability

Browse all Probability questions for SSC CGL, or explore more Quantitative Aptitude topics.

Start a Probability lesson →
Related questions